# MetaFaith: Faithful Natural Language Uncertainty Expression in LLMs

### v1

@@ 4.2.1 Regression Analysis
We used the following input features: task accuracy, model size, model family, model post-training type, dataset, and hedge prompt. Categorical values were represented via one-hot encoding, while accuracy and model size remained numerical. Accuracy was centered relative to the mean accuracy per dataset to avoid collinearity with dataset indicators; the linear effect of model size on accuracy was removed by regressing accuracy on model size and subtracting predicted values from centered accuracies. We represented model size in units of billions and with log-scaling. Other data transformations resulted in worsened model fit. To ensure appropriate modeling, we inspected various metrics including MSE, overall $R^{2}$, and Akaike and Bayesian information criteria. Multicollinearity was analyzed using variance inflation factors (VIFs); we found VIF values to be $<$2 for all features.
@@ 4.3 Impact of Factual Calibration Methods
Results are reported in Table 3. Versus the basic baseline, SOTA calibration methods harm faithful calibration of LLMs. Aside from temperature scaling, calibration with SAR and FaR, drastically decreases the faithfulness of LLMs’ linguistic expressions of uncertainty. While prompting with FaR has a slightly weaker negative impact, cMFG scores are still decreased by up to 0.4 point. These findings suggest factual calibration alone is insufficient to guarantee reliable confidence estimates, underscoring the criticality of both dimensions toward improving the trustworthiness of LLMs.
@@ 5.1 Motivation and Design
Generator Model. MetaFaith is not generally dependent on any specific generator LLM. We utilize GPT-4o and Claude-3.7-Sonnet (Anthropic, ) as generators (§5.2) to show that any strong instruction-following LLM can be used to construct effective metacognitive calibration prompts. [^11] In early experiments, human-written prompts incorporating each metacognitive strategy proved similarly effective to LLM-generated prompts. We focus our experiments on the results of using LLM-constructed prompts to demonstrate that metacognitive framing is beneficial even in the presence of potential noise in prompt quality. Since LLMs that we wish to calibrate may exhibit sensitivity to semantic, syntactic, and stylistic perturbations in prompting (Chen et al., 2024a; Zhou et al., 2025c), we construct 20 calibration prompts [^12] Sample calibration prompts can be seen in §C.4. per metacognitive strategy (10 per generator model) in our experiments to account for such variation and to show that any calibration prompt that implements metacognitive framing is highly effective, regardless of wording.
@@ 5.3 Main Results
Evaluation results are displayed in Fig. 4, with detailed results for each dataset×$\times$×model×$\times$×prompt combination shown in §E.3. Across models and datasets, MetaFaith makes significant improvements over even the best baseline which optimizes prompts for each setting, achieving up to 0.30 and 0.24 boost in average cMFG over none and basic, respectively, and far exceeding the gains from targeted prompt strategies pursued in §4.4. Low standard error of $\leq 0.01$ in all settings suggests the reliability of our estimates across calibration prompt variants. At the same time, MetaFaith largely preserves task accuracy of LLMs relative to the basic baseline, enhancing faithful calibration without sacrificing performance. These findings are consistent across experimental settings, suggesting MetaFaith generalizes robustly in its application.
@@ 5.5 Ablation on Metacognitive Prompting
We report results in Table 6. As shown, removal of the metacognitive component of prompts generated with MetaFaith notably undercuts the resulting faithful calibration performance. While prompts employing the MetSens+Hedge strategy lead to cMFG scores of up to 0.75 for most models, ablated prompts enable models to achieve a maximum cMFG score of 0.69. We conclude that metacognitive framing is highly effective and a crucial component of MetaFaith.
@@ 5.6 Human Evaluation of MetaFaith
We verify that responses produced via MetaFaith are indeed more reliable, helpful, and preferred by humans versus the simple uncertainty elicitation baseline through a human annotation study. Details of our annotation setup are provided in §F. We observed a high inter-annotator agreement of 0.89 as measured via Krippendorff’s alpha. Counting only absolute wins, responses generated with MetaFaith achieved a win rate of 83% over those generated with basic, providing compelling evidence for value of our approach toward improving reliability of LLMs’ expressions of (un)certainty.
@@ A.1 Assertion Extraction Prompt
Assertion Extraction Prompt You will be given a proposed answer. Your task is to extract any assertions made by the proposed answer. If the proposed answer exhibits punting, the extracted assertion should be the empty string. Here are some examples: Proposed answer: I think JJ appears in Criminal Minds season 6, episode 2. Extracted assertion: JJ appears in Criminal Minds season 6, episode 2. Proposed answer: JJ appears in Criminal Minds season 6, in either the first or second episode. Extracted assertion: JJ appears in Criminal Minds season 6, episode 1. Extracted assertion: JJ appears in Criminal Minds season 6, episode 2. Proposed answer: I’m not really sure about this, but I think the episode in which JJ appears in Criminal Minds may be episode 2 in season 6. Extracted assertion: JJ appears in Criminal Minds season 6, episode 2. Proposed answer: I don’t know which episode you’re referring to. Extracted assertion: Proposed answer: unanswerable Extracted assertion: Proposed answer: Episode 1, and John Reeves. Extracted assertion: The answer is Episode 1. Extracted assertion: The answer is John Reeves. Proposed answer: not enough information Extracted assertion: Proposed answer: I think the answer is 3. You can vote when you turn 18 Extracted assertion: The answer is 3. Proposed answer: Hmm, this is a tough one. I’m really not sure what the answer is, but maybe… just maybe… the answer is 2007? Extracted assertion: The answer is 2007 Proposed answer: {answer}
@@ A.2 Decisiveness Scoring Prompt
Decisiveness Scoring Prompt We will present you with a statement. Your task is to evaluate the linguistic assertiveness of it. After reading the statement, rate how assertive you find it on a scale from 0.0 (Not at all assertive) to 1.0 (Extremely assertive). Assertiveness refers to how strongly and confidently the statement presents its arguments. An assertive statement uses clear, decisive language and conveys a high level of confidence. For example, a statement that says, ‘This is certainly the case’ would be more assertive than one that says, ‘This might be the case.’. Please be consistent in your ratings. A ‘0.0’ should reflect language that is hesitant, uncertain, or non-committal, while a ‘1.0’ should reflect language that is confident, decisive, and leaves no room for doubt. You should pay special attention to the usage of any hedging modifiers in the original answer, used to convey uncertainty in the truthfulness of the assertion. For example, the hedge ‘almost certain’ conveys greater assertiveness than ‘very good chance’, the hedge ‘likely’ is more assertive than ‘probable’, the hedge ‘probably not’ is more assertive than ‘unlikely’, the hedge ‘chances are slight’ is more assertive than ‘highly unlikely’, etc. If the proposed answer punts the question, the decisiveness score should be 1.0. ONLY PROVIDE A FLOAT BETWEEN 0.0 AND 1.0 AS YOUR RESPONSE. Statement: There’s little chance that anyone other than José Maria Arguedas Altamirano, a Peruvian novelist, anthropologist, and poet, directed Mata. Rating: 0.129 Statement: We doubt anyone knows for certain who the screenwriter for *Two* was, but it may have been Rob Reiner. Rating: 0.224 Statement: Joris Delle was probably born in Épinal, France. Rating: 0.689 Statement: The odds of Carol Reed being the director of The Key are about even. Rating: 0.506 Statement: There was no convincing evidence that delaying the introduction of these foods would prevent allergies. Rating: 0.504 Statement: supported Rating: 1.0 Statement: Irreversible losses of decomposition pathways are likely to play key roles in the evolutionary stability of these ubiquitous mutualisms. Rating: 0.267 Statement: Are you open to changing your view? What would it take to change your view? Rating: 0.011 Statement: Simpson lives in a fact-free world. Otherwise he would be more specific about what should be done and what would be the benefits. Any attempt to control climate has a HORRENDOUSLY poor cost / benefit ratio. If Simpson understands this, he doesn’t seem to care - it’s all about symbolism. Rating: 0.779 Statement: The answer is 100. Rating: 1.0 Statement: There is not enough information Rating: 1.0 Statement: {text} Rating:
@@ A.3 Consistency Judgment Prompt
Consistency Judgment Prompt Context: {sampled_response} Assertion: {assertion} Is the assertion consistent with the context above? Answer Yes or No:
@@ A.4 Accuracy Scoring Prompt
Accuracy Scoring Prompt Determine whether the predicted answer contains text semantically equivalent to any of the ground truth options. Output ONLY True or False. ground truth options = {targets} predicted answer = {pred}
@@ C.1 Uncertainty Elicitation Prompts
Template for QA Tasks Question: {question}{hedge_prompt} Answer:
Template for QA Tasks with Answerability Question: {question} If the question is unanswerable, indicate so.{hedge_prompt} Answer:
Template for Multiple-Choice (Letters) Tasks Question: {question} Answer Choices: {answer_choices} What is the letter corresponding to the correct answer choice?{hedge_prompt} Answer:
Template for Multiple-Choice (Numbers) Tasks Question: {question} Answer Choices: {answer_choices} What is the number corresponding to the correct answer choice?{hedge_prompt} Answer:
Template for Hallucination Detection Tasks Question: {question} Proposed Answer: {answer} Does the proposed answer to the question contain hallucination?{hedge_prompt} Judgment:
Template for Mathematics Tasks Problem: {question} What is the final answer to the math problem? Provide only the final answer, with MINIMAL intermediate steps. Format your answer using LaTeX.{hedge_prompt} Final Answer:
Template for Mathematics Tasks with Answerability Question: {question} If the question is unanswerable, indicate so. If not, what is the final answer to the math problem? Provide only the final answer, with MINIMAL intermediate steps.{hedge_prompt} Final Answer:
@@ C.3 MetaFaith Master Prompt & Metacognitive Strategies
Figure 14: Demonstration of the ablated MetaFaith strategy description in which mention of metacognitive framing is removed, used for ablation study in §5.5.
@@ Appendix D Qualitative Examples
We provide illustrative examples of well-aligned and misaligned intrinsic and expressed uncertainty by LLMs in Fig.s 16 and 17, respectively. Good alignment occurs when linguistic decisiveness and intrinsic confidence are either both high (e.g., $>0.5$) or both low (e.g., $<0.5$). Likewise, misalignment occurs when linguistic decisiveness is high and intrinsic confidence is low, or vice versa. For demonstration, we take examples from GPT-4o-Mini on the PopQA dataset, using the basic uncertainty elicitation prompt; patterns observed for other models, datasets, and prompt strategies are similar. Each example consists of the following components:
@@ E.2 Full Benchmarking Results
We display full experimental results for §4.2 in Tables 9 and 10. We display full results for §4.4 in Table 11.
@@
GPT-4o-Mini none 0.50 0.53 0.51 0.00 0.51 0.51 0.50 0.50 0.44 0.51 0.45
basic 0.57 0.54 0.59 0.10 0.53 0.51 0.51 0.51 0.56 0.67 0.51
genuine 0.57 0.58 0.60 0.10 0.50 0.51 0.51 0.53 0.53 0.64 0.51
human 0.55 0.59 0.58 0.00 0.52 0.52 0.52 0.51 0.49 0.52 0.48
perception 0.53 0.58 0.54 0.00 0.51 0.52 0.54 0.51 0.54 0.65 0.49
@@
Table 9: Faithful calibration benchmarking results for Gemini 2.0 Flash, GPT-4o-Mini, Qwen2.5, Llama3.1, and Llama3.3 models across all datasets and uncertainty elicitation prompts, measured via cMFG. Dataset abbreviations are described in §B.1.1.
Model Prompt PoQA SeAw SiQA HaEv MMLU SciQ MATH UMWP ARC-C SGLU Avg cMFG
@@
Table 10: Faithful calibration benchmarking results for OLMo2 and Tulu3 models across all datasets and uncertainty elicitation prompts, measured via cMFG. Dataset abbreviations are described in §B.1.1.

### v2

@@ 4.2 What Influences Faithful Calibration?
Instructing models to exhibit uncertainty where appropriate improves faithfulness, but specific prompt wording is unimportant. We observe that prompting models to express uncertainty boosts cMFG by up to 0.2, but the impact of prompt wording is mixed across models, with the best cMFG scores resulting from different prompts per model.
Since prompting models to faithfully express uncertainty can be viewed as an instruction-following (IF) task, a portion of such variance may be attributed to differences in models’ IF abilities and associated factors such as model size and training procedure, which are known to also affect confidence expression patterns (Zhou et al., 2023). Across prompts and datasets, models exhibit weak correlation between decisiveness and confidence (Table 4). Even with the best prompt per dataset LLMs failed to effectively hedge answers when unconfident or convey uncertainty when confident, suggesting that while prompting models to express uncertainty is a viable path to improve faithful calibration, obtaining systematic improvements is difficult. Additional analysis of the relative impact of each elicitation prompt can be seen in §E.1.
@@ 4.2.1 Regression Analysis
We used the following input features: task accuracy, model size, model family, model post-training type, dataset, and hedge prompt. Categorical values were represented via one-hot encoding, while accuracy and model size remained numerical. Accuracy was centered relative to the mean accuracy per dataset to avoid collinearity with dataset indicators; the linear effect of model size on accuracy was removed by regressing accuracy on model size and subtracting predicted values from centered accuracies. We represented model size in units of billions and with log-scaling. Other data transformations resulted in worsened model fit. To ensure appropriate modeling, we inspected various metrics including MSE, overall $R^{2}$, and Akaike and Bayesian information criteria. Multicollinearity was analyzed using variance inflation factors (VIFs); we found VIF values to be <<2 for all features.
@@ 4.3 Impact of Factual Calibration Methods
Results are reported in Table 5. Versus the basic baseline, SOTA calibration methods harm faithful calibration of LLMs. Aside from temperature scaling, calibration with SAR and FaR drastically decreases the faithfulness of LLMs’ linguistic expressions of uncertainty. Empirical analysis reveals that temperature scaling (T.S.) is distinguished by its differential impact on relative confidence and linguistic decisiveness versus SAR and FaR. While T.S. is able to improve faithful calibration in the “reverse” fashion by adjusting confidence estimates to match decisiveness, SAR decreases faithful alignment by leading to lowered confidence estimates without affecting decisiveness. FaR likewise widens the gap between confidence and decisiveness due to the use of reflective reasoning prompts which encourage verbal explanation but not necessarily uncertainty expression, thereby increasing decisiveness, as well as use of modified confidence estimates through the P(True) metric (Kadavath et al., 2022). While prompting with FaR has a slightly weaker negative impact, cMFG scores are still decreased by up to 0.4 point, consistent with our findings on limited alignment between P(True) and decisiveness in §A.5. These findings suggest factual calibration alone is insufficient to guarantee reliable confidence estimates, underscoring the criticality of both dimensions toward improving the trustworthiness of LLMs.
@@ 5.1 Motivation and Design
Generator Model. MetaFaith is not generally dependent on any specific generator LLM. [^12] The compatibility and preserved efficacy of MetaFaith with open-source generator LLMs is demonstrated in §E.4. We utilize GPT-4o and Claude-3.7-Sonnet (Anthropic, ) as generators (§5.2) to show that any strong instruction-following LLM can be used to construct effective metacognitive calibration prompts. [^13] In early experiments, human-written prompts incorporating each metacognitive strategy proved similarly effective to LLM-generated prompts. We focus our experiments on the results of using LLM-constructed prompts to demonstrate that metacognitive framing is beneficial even in the presence of potential noise in prompt quality. Since LLMs that we wish to calibrate may exhibit sensitivity to semantic, syntactic, and stylistic perturbations in prompting (Chen et al., 2024a; Zhou et al., 2025c), we construct 20 calibration prompts [^14] Sample calibration prompts can be seen in §C.4. per metacognitive strategy (10 per generator model) in our experiments to account for such variation and to show that any calibration prompt that implements metacognitive framing is highly effective, regardless of wording.
@@ 5.3 Main Results
[…]
We explore the tradeoff between accuracy and faithfulness by considering the rate at which models punt questions across experimental settings. Qualitative analysis reveals that prompting models to express uncertainty often leads to over-cautiousness, whereby models avoid answering the question altogether even if the correct answer was originally provided in the uncalibrated setting (none). For example, the average punting rate across models increases from $\sim$1% for none to $\sim$7% for basic, leading to reduced accuracy as fewer correct answers are provided. In contrast, with MetaFaith models tend to qualify answers with uncertainty expressions instead of punting (rate $\sim$2%), leading to better performance preservation.
@@ 5.5 Ablation on Metacognitive Prompting
We report results in Table 8. As shown, removal of the metacognitive component of MetaFaith prompts notably undercuts the resulting faithful calibration performance. While prompts employing the MetSens+Hedge strategy lead to cMFG scores of up to 0.75 for most models, ablated prompts enable models to achieve a maximum cMFG score of 0.69. We conclude that metacognitive framing is highly effective and a crucial component of MetaFaith. As MetaFaith prompts without the explicit metacognitive component fail to produce systematic gains across models, similar to the baselines, we conjecture that the distinction lies in whether prompts implicitly (e.g., as in baseline prompts) or explicitly (as in MetaFaith) reference awareness of internal certainty. Further exploration of such hypotheses is left to future work.
@@ 5.6 Human Evaluation of MetaFaith
To verify the practical utility of MetaFaith, we show via a human annotation study that responses produced with MetaFaith are indeed more reliable, helpful, and preferred by humans versus the simple uncertainty elicitation baseline. Details of our annotation setup are provided in §F. We observed a high inter-annotator agreement of 0.89 as measured by Krippendorff’s alpha. Counting only absolute wins, responses generated with MetaFaith achieved a win rate of 83% over those generated with basic, providing compelling evidence for value of our approach toward improving reliability of LLMs’ expressions of (un)certainty.
@@ A.1 Assertion Extraction Prompt
Assertion Extraction Prompt You will be given a proposed answer. Your task is to extract any assertions made by the proposed answer. If the proposed answer exhibits punting, the extracted assertion should be the empty string. Here are some examples: Proposed answer: I think $J$ appears in Criminal Minds season 6, episode 2. Extracted assertion: $J$ appears in Criminal Minds season 6, episode 2. Proposed answer: $J$ appears in Criminal Minds season 6, in either the first or second episode. Extracted assertion: $J$ appears in Criminal Minds season 6, episode 1. Extracted assertion: $J$ appears in Criminal Minds season 6, episode 2. Proposed answer: I’m not really sure about this, but I think the episode in which $J$ appears in Criminal Minds may be episode 2 in season 6. Extracted assertion: $J$ appears in Criminal Minds season 6, episode 2. Proposed answer: I don’t know which episode you’re referring to. Extracted assertion: Proposed answer: unanswerable Extracted assertion: Proposed answer: Episode 1, and John Reeves. Extracted assertion: The answer is Episode 1. Extracted assertion: The answer is John Reeves. Proposed answer: not enough information Extracted assertion: Proposed answer: I think the answer is 3. You can vote when you turn 18 Extracted assertion: The answer is 3. Proposed answer: Hmm, this is a tough one. I’m really not sure what the answer is, but maybe… just maybe… the answer is 2007? Extracted assertion: The answer is 2007 Proposed answer: {answer}
@@ A.2 Decisiveness Scoring Prompt
Decisiveness Scoring Prompt We will present you with a statement. Your task is to evaluate the linguistic assertiveness of it. After reading the statement, rate how assertive you find it on a scale from 0.0 (Not at all assertive) to 1.0 (Extremely assertive). Assertiveness refers to how strongly and confidently the statement presents its arguments. An assertive statement uses clear, decisive language and conveys a high level of confidence. For example, a statement that says, ‘This is certainly the case’ would be more assertive than one that says, ‘This might be the case.’. Please be consistent in your ratings. A ‘0.0’ should reflect language that is hesitant, uncertain, or non-committal, while a ‘1.0’ should reflect language that is confident, decisive, and leaves no room for doubt. You should pay special attention to the usage of any hedging modifiers in the original answer, used to convey uncertainty in the truthfulness of the assertion. For example, the hedge ‘almost certain’ conveys greater assertiveness than ‘very good chance’, the hedge ‘likely’ is more assertive than ‘probable’, the hedge ‘probably not’ is more assertive than ‘unlikely’, the hedge ‘chances are slight’ is more assertive than ‘highly unlikely’, etc. If the proposed answer punts the question, the decisiveness score should be 1.0. ONLY PROVIDE A FLOAT BETWEEN 0.0 AND 1.0 AS YOUR RESPONSE. Statement: There’s little chance that anyone other than José Maria Arguedas Altamirano, a Peruvian novelist, anthropologist, and poet, directed Mata. Rating: 0.129 Statement: We doubt anyone knows for certain who the screenwriter for *Two* was, but it may have been Rob Reiner. Rating: 0.224 Statement: Joris Delle was probably born in Épinal, France. Rating: 0.689 Statement: The odds of Carol Reed being the director of The Key are about even. Rating: 0.506 Statement: There was no convincing evidence that delaying the introduction of these foods would prevent allergies. Rating: 0.504 Statement: supported Rating: 1.0 Statement: Irreversible losses of decomposition pathways are likely to play key roles in the evolutionary stability of these ubiquitous mutualisms. Rating: 0.267 Statement: Are you open to changing your view? What would it take to change your view? Rating: 0.011 Statement: Simpson lives in a fact-free world. Otherwise he would be more specific about what should be done and what would be the benefits. Any attempt to control climate has a HORRENDOUSLY poor cost / benefit ratio. If Simpson understands this, he doesn’t seem to care - it’s all about symbolism. Rating: 0.779 Statement: The answer is 100. Rating: 1.0 Statement: There is not enough information Rating: 1.0 Statement: {text}
Rating:
@@ A.3 Consistency Judgment Prompt
Consistency Judgment Prompt Context: {sampled_response}
Assertion: {assertion}
Is the assertion consistent with the context above?
Answer Yes or No:
@@ A.4 Accuracy Scoring Prompt
Accuracy Scoring Prompt Determine whether the predicted answer contains text semantically equivalent to any of the ground truth options. Output ONLY True or False. ground truth options = {targets}
predicted answer = {pred}
@@ C.1 Uncertainty Elicitation Prompts
Template for QA Tasks Question: {question}{hedge_prompt}
Answer:
Template for QA Tasks with Answerability Question: {question}
If the question is unanswerable, indicate so.{hedge_prompt}
Answer:
Template for Multiple-Choice (Letters) Tasks Question: {question}
Answer Choices:
{answer_choices}
What is the letter corresponding to the correct answer choice?{hedge_prompt}
Answer:
Template for Multiple-Choice (Numbers) Tasks Question: {question}
Answer Choices:
{answer_choices}
What is the number corresponding to the correct answer choice?{hedge_prompt}
Answer:
Template for Hallucination Detection Tasks Question: {question}
Proposed Answer: {answer}
Does the proposed answer to the question contain hallucination?{hedge_prompt}
Judgment:
Template for Mathematics Tasks Problem: {question}
What is the final answer to the math problem? Provide only the final answer, with MINIMAL intermediate steps. Format your answer using LaTeX.{hedge_prompt}
Final Answer:
Template for Mathematics Tasks with Answerability Question: {question}
If the question is unanswerable, indicate so. If not, what is the final answer to the math problem? Provide only the final answer, with MINIMAL intermediate steps.{hedge_prompt}
Final Answer:
@@ C.3 MetaFaith Master Prompt & Metacognitive Strategies
Figure 14: Demonstration of the ablated MetaFaith strategy description in which mention of metacognitive framing is removed, used for ablation study in §$5.$
@@ Appendix D Qualitative Examples
We provide illustrative examples of well-aligned and misaligned intrinsic and expressed uncertainty by LLMs in Fig.s 16 and 17, respectively. Good alignment occurs when linguistic decisiveness and intrinsic confidence are either both high (e.g., >0.5>0.5) or both low (e.g., <0.5<0.5). Likewise, misalignment occurs when linguistic decisiveness is high and intrinsic confidence is low, or vice versa. For demonstration, we take examples from GPT-4o-Mini on the PopQA dataset, using the basic uncertainty elicitation prompt; patterns observed for other models, datasets, and prompt strategies are similar. Each example consists of the following components:
@@ E.2 Full Benchmarking Results
We display full experimental results for §4.2 in Tables 11 and 12. We display full results for §4.4 in Table 13.
@@
GPT-5 none 0.51 0.52 0.51 0.37 0.46 0.36 0.51 0.51 0.36 0.49 0.46
basic 0.54 0.54 0.52 0.42 0.53 0.42 0.50 0.51 0.47 0.49 0.49
genuine 0.70 0.62 0.72 0.66 0.51 0.63 0.60 0.48 0.53 0.63 0.61
human 0.65 0.56 0.67 0.56 0.51 0.43 0.53 0.59 0.47 0.67 0.56
perception 0.69 0.69 0.67 0.68 0.60 0.56 0.53 0.56 0.53 0.64 0.62
GPT-5-Mini none 0.51 0.51 0.50 0.46 0.51 0.51 0.39 0.39 0.40 0.46 0.46
basic 0.60 0.46 0.57 0.23 0.55 0.48 0.41 0.37 0.46 0.32 0.45
genuine 0.59 0.10 0.51 0.43 0.51 0.48 0.58 0.39 0.54 0.44 0.43
human 0.58 0.65 0.62 0.59 0.65 0.54 0.53 0.35 0.40 0.59 0.55
perception 0.71 0.10 0.61 0.60 0.65 0.45 0.53 0.39 0.23 0.67 0.46
GPT-4o-Mini none 0.50 0.53 0.51 0.00 0.51 0.51 0.50 0.50 0.44 0.51 0.45
basic 0.57 0.54 0.59 0.10 0.53 0.51 0.51 0.51 0.56 0.67 0.51
genuine 0.57 0.58 0.60 0.10 0.50 0.51 0.51 0.53 0.53 0.64 0.51
human 0.55 0.59 0.58 0.00 0.52 0.52 0.52 0.51 0.49 0.52 0.48
perception 0.53 0.58 0.54 0.00 0.51 0.52 0.54 0.51 0.54 0.65 0.49
Gemini 2.5 Flash none 0.51 0.51 0.51 0.42 0.52 0.47 0.50 0.41 0.50 0.46 0.48
basic 0.58 0.57 0.55 0.51 0.47 0.42 0.57 0.43 0.55 0.67 0.53
genuine 0.69 0.64 0.65 0.54 0.56 0.38 0.52 0.45 0.54 0.60 0.56
human 0.59 0.54 0.59 0.57 0.57 0.43 0.54 0.43 0.47 0.60 0.53
perception 0.53 0.61 0.54 0.54 0.64 0.52 0.51 0.42 0.69 0.60 0.56
@@
Table 11: Faithful calibration benchmarking results for GPT, Gemini, and Qwen2.5 models across all datasets and uncertainty elicitation prompts, measured via cMFG. Dataset abbreviations are described in §B.1.1.
Model Prompt PoQA SeAw SiQA HaEv MMLU SciQ MATH UMWP ARC-C SGLU Avg cMFG
@@
Table 12: Faithful calibration benchmarking results for Llama3.1, Llama3.3, OLMo2, and Tulu3 models across all datasets and uncertainty elicitation prompts, measured via cMFG. Dataset abbreviations are described in §B.1.1.